# Zhan Gaoa, Ji Hyung Leeb, Ziwei Meic, Zhentao Shic

### v2

@@ 4.1 Setup
We consider the linear predictive regression model in (2.1). The vector of stationary components, denoted as $v_{t}=\left(u_{t},\Delta X_{t}^{\top},Z_{t}^{\top}\right)^{\top}$, is generated by the two different processes:
@@
where $R_{n}=\text{diag}(0,0.3,0.3,\ldots,0.3)$. Under this $R_{n}$ the error term $u_{t}$ in the main regression (2.1) remains i.i.d. satisfying the m.d.s. condition in Assumption 1, while the unit root generators $\Delta X_{t}$ and the stationary regressors $Z_{t}$ are AR(1) processes. The covariance matrix $\Sigma=(\Sigma_{ij})_{i,j\in[p]}$ is specified as
@@
Recall $0_{p}$ is a $p$-dimensional zero vector and $1_{p}$ is a $p$-dimensional vector of ones. The true coefficient vectors are:
$\beta^{*}=\left(\beta_{1}^{*},\dfrac{0.5}{\sqrt{n}}\times 1_{4}^{\top},0_{p_{x}-5}^{\top}\right)^{\top},\ \ \gamma^{*}=(\gamma_{1}^{*},0.5\times 1_{2}^{\top},0.25\times 1_{2}^{\top},0_{p_{z}-5}^{\top})^{\top}.$
The specification involves four active unit root regressors and four active stationary regressors. The $1/\sqrt{n}$ scaling balances the regression by normalizing the coefficients of the unit root regressors. We test the hypothesis $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$ with sample sizes $n\in\{200,300,400,500,600\}$ and the dimensionality pairs $(p_{x},p_{z})\in\{(50,100),(100,150),(150,300)\}.$ We conduct 2000 replications for each setting.
We compare the finite sample performance of the testing procedure XDlasso, as described in Algorithm 1, and Dlasso with the score vector in (2.14) and (2.15). In addition, we consider two infeasible testing procedures as benchmarks. Using the known active set of regressors, we conduct IVX inference (IVX oracle) and the standard $t$-test based on the OLS estimator (OLS oracle), employing only the regressor of interest and the active regressors, which form a low-dimensional predictive regression model. We set $C_{\zeta}=5$ and $\tau=0.5$ for the parameter $\rho_{\zeta}$ specified as (3.5). As discussed in Remark 5, the choice $\tau=$ 0.5 admits the weakest sparsity condition, and thus effectively improves the finite sample performance.
Both XDlasso and Dlasso involve Slasso, where the selection of tuning parameters $\lambda$ and $\mu$ affects finite sample performance. In our experiments, we employ the block 10-fold cross-validation (CV), splitting the sample into 10 equally sized chronologically ordered consecutive blocks for validations. In Theorem 1, the tuning parameters are specified as constants multiplied by the appropriate rates of convergence determined by the sample size $n$, dimensionality $p$ and the mixing condition constant $r$. As a benchmark, we also calibrate the tuning parameters following Lee et al., (2022) to examine the validity of the theoretical orders of tuning parameters specified in Theorem 1. Specifically, we perform $500$ pilot replications for each DGP, with $n_{0}=400$ , $\left(p_{x0},p_{z0}\right)=\left(100,150\right)$, and $p_{0}=p_{x0}+p_{z0}$. In each replication $q=1,2,\dots,500$, we use the 10-fold cross-validation to choose the tuning parameters $\lambda^{(q)}$ and $\mu^{(q)}$, and calibrate the constants as
@@ 4.2 Results
We first investigate the empirical size of different testing methods at a 5% nominal significance level when the true coefficients $\beta_{1}^{*}=0$ and $\gamma_{1}^{*}=0$. Tables 1 and 2 report the empirical sizes under the IID and AR(1) innovations, respectively. Foremost among the findings is that XDlasso effectively controls the empirical size for both $\beta_{1}^{*}$, associated with a unit root regressor, and $\gamma_{1}^{*}$, associated with a stationary regressor. This performance stands in sharp contrast to that of Dlasso and OLS oracle, which exhibit severe size distortions for $\beta_{1}^{*}$. Such distortions can be attributed to the failure to account for the Stambaugh bias arising from nonstationarity. Furthermore, the results yield noteworthy insights regarding the tuning parameter selection. The empirical size of XDlasso with both cross-validated and calibrated tuning parameters is aligned with the nominal level. This result not only validates the asymptotic rates of tuning parameters specified in Theorem 1 but also supports CV as a feasible data-driven tuning parameter selection method in practice. Lastly, we investigate the efficiency and robustness of XDlasso in comparison with alternatives. When compared to the unbiased but infeasible “IVX oracle” estimator, the confidence intervals produced by XDlasso are only slightly wider. In contrast to the estimators solely for low-dimensional data, XDlasso demonstrates robustness by accommodating high dimensional covariates without compromising much in efficiency.
@@ IVX Oracle OLS Oracle XDlasso Dlasso XDlasso Dlasso
200 0.051 0.365 0.061 0.320 0.073 0.331 0.077 0.294 0.078 0.319 0.080 0.284
300 0.041 0.290 0.057 0.262 0.072 0.266 0.067 0.241 0.073 0.261 0.070 0.237
400 0.046 0.247 0.061 0.226 0.072 0.229 0.064 0.210 0.067 0.227 0.065 0.209
500 0.055 0.219 0.060 0.203 0.065 0.204 0.066 0.189 0.064 0.203 0.067 0.188
600 0.047 0.198 0.058 0.185 0.062 0.185 0.065 0.173 0.063 0.185 0.063 0.173
Notes: The data generating process corresponds to (4.1). The upper and lower panels report the empirical size of testing the null hypotheses $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$, respectively, at a 5% nominal significance level. “Size” is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (2.13) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}$ is the 97.5-th percentile of the standard normal distribution. “Len.” refers to the median length of the 95% confidence intervals across replications. The IVX oracle and OLS oracle are infeasible estimators. The “Calibrated” and “CV” columns refer to the methods used for choosing the tuning parameters through calibration and cross-validation, respectively.
@@
Notes: The data generating process corresponds to (4.1). The upper and lower panels report the empirical size of testing the null hypotheses $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$ at a 5% nominal significance level, respectively. “Size” is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (2.13) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}$ is the 97.5-th percentile of the standard normal distribution. “Len.” refers to the median length of the 95% confidence intervals across replications. The IVX oracle and OLS oracle are infeasible estimators. The “Calibrated” and “CV” columns refer to the methods used for choosing the tuning parameters through calibration and cross-validation, respectively.
@@
We now turn to the empirical power of the XDlasso inference. Figure 2 plots power curves for the null hypotheses $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$ under varying true coefficient values. In our analysis, we vary the coefficient $\beta_{1}^{*}$ or $\gamma_{1}^{*}$ from 00 to $0.5$ while maintaining the other equal to 00. Across various configurations, XDlasso exhibits increasingly high power against the null hypothesis as the true coefficient moves away from 0 and the sample size $n$ increases. Furthermore, Figure 3 compares the power for testing $\beta_{1}^{*}$ and $\gamma_{1}^{*}$, which reveals that the power associated with a unit root regressor surpasses that of a stationary regressor. This observation provides finite sample evidence to support the theoretical results in Theorem 2. Specifically, it corroborates the faster convergence of standard error for unit root regressors compared with stationary regressors, thereby inducing higher power in the hypothesis testing.
@@
Notes: The left and right panels correspond to DGPs (4.1) and (4.2), respectively. In each subplot, the first row depicts to the empirical power function for $\beta_{1}^{*}$, associated with a nonstationary regressor, across various $\left(p_{x},p_{z}\right)$ configurations, while the second row pertains to $\gamma_{1}^{*}$, associated with a stationary regressor. The empirical power is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (2.13) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}$ is the 97.5-th percentile of the standard normal distribution.
@@
Notes: This figure plots the power curves under $n=600.$ The first and second rows correspond to DGPs (4.2) and (4.1), respectively. In each subplot, blue lines represent the empirical power function for $\beta_{1}^{*}$, associated with a nonstationary regressor, while red lines represent that for $\gamma_{1}^{*}$, associated with a stationary regressor. Empirical power is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (2.13) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}$ is the 97.5-th percentile of the standard normal distribution.
@@ 5 Empirical Applications
In recent years, high dimensional macroeconomic data have been extensively used to forecast key variables of interest; see Smeekes and Wijler, (2018), Medeiros et al., (2021), and Giannone et al., (2021), for example. Researchers have primarily focused on the point estimation of forecast. There has been limited empirical exploration of statistical inference on predictive power of specific predictors, due to a lack of suitable toolkits. This section showcases two empirical applications using our proposed XDlasso inference method. We utilize the monthly data of 112 U.S. macroeconomic variables spanning from January 1960 to April 2024, sourced from the FRED-MD dataset by McCracken and Ng, (2016).

### v3

@@ 4.1 Setup
We consider the linear predictive regression model in (2.1). The vector of the stationary components, denoted as $v_{t}=\left(u_{t},e_{t}^{\top},Z_{t}^{\top}\right)^{\top}$, is generated by the two different processes:
@@
where $R_{n}=\text{diag}(0,0.3,0.3,\ldots,0.3)$. Under this $R_{n}$ the error term $u_{t}$ in the main regression (2.1) remains i.i.d., satisfying the m.d.s. condition in Assumption 1, while the (local) unit root innovations $e_{t}$ and the stationary regressors $Z_{t}$ are AR(1) processes. The covariance matrix $\Sigma=(\Sigma_{ij})_{i,j\in[p]}$ is specified as
@@
The persistent regressors $X_{t}$ are generated by
$X_{t}=\text{diag}(\rho^{*})X_{t-1}+e_{t},$ (4.3)
where $\rho^{*}=(1,1-1/n,1+1/n,1,1-1/n,1+1/n,\cdots)^{\top}\in\mathbb{R}^{p_{x}}$. Recall $0_{p}$ is a $p$-dimensional zero vector and $1_{p}$ is a $p$-dimensional vector of ones. The true coefficient vectors are:
$\beta^{*}=\left(\beta_{1}^{*},\dfrac{0.5}{\sqrt{n}}\times 1_{4}^{\top},0_{p_{x}-5}^{\top}\right)^{\top},\ \ \gamma^{*}=(\gamma_{1}^{*},0.5\times 1_{2}^{\top},0.25\times 1_{2}^{\top},0_{p_{z}-5}^{\top})^{\top}.$ (4.4)
The specification involves four active LUR regressors and four active stationary regressors. [^3] In Section B.1, we consider a setting with more nonzero coefficients and find robust performance of XDlasso. The $1/\sqrt{n}$ scaling balances the regression by normalizing the coefficients of the unit root regressors. We test the hypothesis $\mathbb{H}_{0}:\beta_{1}^{*}=0$, $\mathbb{H}_{0}:\gamma_{1}^{*}=0$, and the joint null hypothesis $\mathbb{H}_{0}:\beta_{1}^{*}=\gamma_{1}^{*}=0$, respectively, under the sample sizes $n\in\{200,300,400,500,600\}$ and the dimensionality pairs $(p_{x},p_{z})\in\{(50,100),(100,150),(150,300)\}.$ We conduct 2000 replications in each setting.
We compare the finite sample performance of XDlasso, as described in Algorithm 1, and Dlasso with the score vector in (2.16) and (2.17). In addition, we consider two infeasible testing procedures as benchmarks. Using the known active set of regressors, we conduct IVX inference (IVX oracle) and the standard $t$-test based on the OLS estimator (OLS oracle), employing only the regressor of interest and the active regressors, which form a low-dimensional predictive regression model. We set $C_{\zeta}=5$ and $\tau=0.5$ for the parameter $\rho_{\zeta}$ specified as (3.5). As discussed in Remark 8, the choice $\tau=$ 0.5 admits the weakest sparsity condition, and thus effectively improves the finite sample performance.
Both XDlasso and Dlasso involve Slasso, where the selection of tuning parameters $\lambda$ and $\mu$ affects finite sample performance. In our experiments, we employ the block 10-fold cross-validation (CV), splitting the sample into 10 equally sized chronologically ordered consecutive blocks for validations. Though the unconditional variances of nonstationary regressors vary in different chronological blocks, the standardization in the Slasso estimators like (2.4) and (2.12) account for such variation. This explains the robustness of the block CV to nonstationary time series, as shown by our simulation results.
In Theorem 1, the tuning parameters are specified as constants multiplied by the appropriate rates of convergence determined by the sample size $n$, dimensionality $p$ and the mixing condition constant $r$. As a benchmark, we also calibrate the tuning parameters following lee2022lasso to examine the validity of the theoretical orders of tuning parameters specified in Theorem 1. Specifically, we perform $500$ pilot replications for each DGP, with $n_{0}=400$ , $\left(p_{x0},p_{z0}\right)=\left(100,150\right)$, and $p_{0}=p_{x0}+p_{z0}$. In each replication $q=1,2,\dots,500$, we use the 10-fold cross-validation to choose the tuning parameters $\lambda^{(q)}$ and $\mu^{(q)}$, and calibrate the constants as
@@ 4.2 Results
We first investigate the empirical size of different testing methods at a 5% nominal significance level when the true coefficients $\beta_{1}^{*}=0$ and $\gamma_{1}^{*}=0$. Tables 1 and 2 report the empirical sizes under the IID and AR(1) innovations, respectively. Foremost among the findings is that XDlasso effectively controls the empirical size for both $\beta_{1}^{*}$, associated with a unit root regressor, and $\gamma_{1}^{*}$, associated with a stationary regressor. This performance stands in sharp contrast to that of Dlasso and OLS oracle, which exhibit severe size distortions for $\beta_{1}^{*}$. Such distortions can be attributed to the failure to account for the Stambaugh bias arising from nonstationarity. Furthermore, the results yield noteworthy insights regarding the tuning parameter selection. The empirical size of XDlasso with both cross-validated and calibrated tuning parameters is close to the nominal level. This result not only validates the asymptotic rates of tuning parameters specified in Theorem 1 but also supports CV as a feasible data-driven tuning parameter selection method in practice. In addition, we investigate the efficiency and robustness of XDlasso in comparison with alternatives. When compared to the unbiased but infeasible “IVX oracle” estimator, the confidence intervals produced by XDlasso are only slightly wider. In contrast to the estimators solely for low-dimensional data, XDlasso demonstrates robustness by accommodating high dimensional covariates without compromising much in efficiency. Lastly, the empirical sizes for testing the joint null hypothesis $\mathbb{H}_{0}:\beta_{1}^{*}=\gamma_{1}^{*}=0$, as reported in Table 3, are also well controlled around the nominal level across setups. The results validate the theoretical result in Theorem 3 and is consistent with the findings for testing $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$ individually.
@@ IVX Oracle OLS Oracle XDlasso Dlasso XDlasso Dlasso
200 0.040 0.375 0.056 0.324 0.060 0.329 0.056 0.291 0.066 0.317 0.065 0.282
300 0.037 0.297 0.040 0.264 0.054 0.265 0.055 0.241 0.059 0.260 0.056 0.237
400 0.036 0.251 0.046 0.228 0.053 0.228 0.047 0.210 0.053 0.227 0.050 0.208
500 0.030 0.222 0.048 0.204 0.051 0.203 0.052 0.188 0.050 0.203 0.049 0.188
600 0.043 0.201 0.049 0.186 0.057 0.185 0.057 0.172 0.052 0.185 0.054 0.173
Notes: The data generating process corresponds to (4.1). The upper and lower panels report the empirical size of testing the null hypotheses $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$, respectively, at a 5% nominal significance level. “Size” is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (2.15) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}$ ($\approx 1.96$) is the 97.5-th percentile of the standard normal distribution. “Len.” refers to the median length of the 95% confidence intervals across replications. The IVX oracle and OLS oracle are infeasible estimators. The “Calibrated” and “CV” columns refer to the methods used for choosing the tuning parameters through calibration and cross-validation, respectively.
@@
Notes: The data generating process corresponds to (4.2). The upper and lower panels report the empirical size of testing the null hypotheses $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$ at a 5% nominal significance level, respectively. “Size” is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (2.15) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}$ ($\approx$ 1.96) is the 97.5-th percentile of the standard normal distribution. “Len.” refers to the median length of the 95% confidence intervals across replications. The IVX oracle and OLS oracle are infeasible estimators. The “Calibrated” and “CV” columns refer to the methods used for choosing the tuning parameters through calibration and cross-validation, respectively.
Table 3: Empirical size: Joint null hypothesis $\mathbb{H}_{0}:\beta_{1}^{*}=\gamma_{1}^{*}=0$
Case I: IID Innovations
$\left(p_{x},p_{z}\right)=\left(50,100\right)$ $\left(p_{x},p_{z}\right)=\left(100,150\right)$ $\left(p_{x},p_{z}\right)=\left(150,300\right)$
$n$ Calibrated CV Calibrated CV Calibrated CV
200 0.064 0.077 0.067 0.076 0.053 0.069
300 0.057 0.066 0.064 0.081 0.052 0.063
400 0.058 0.069 0.067 0.080 0.047 0.056
500 0.054 0.064 0.061 0.069 0.053 0.063
600 0.051 0.055 0.060 0.069 0.056 0.059
Case II: AR(1) Innovations
$\left(p_{x},p_{z}\right)=\left(50,100\right)$ $\left(p_{x},p_{z}\right)=\left(100,150\right)$ $\left(p_{x},p_{z}\right)=\left(150,300\right)$
$n$ Calibrated CV Calibrated CV Calibrated CV
200 0.070 0.090 0.064 0.098 0.061 0.084
300 0.070 0.095 0.065 0.096 0.055 0.079
400 0.056 0.075 0.063 0.090 0.054 0.070
500 0.059 0.079 0.064 0.086 0.055 0.067
600 0.053 0.071 0.067 0.079 0.060 0.076
Notes: The data generating processes for Case I and Case II correspond to (4.1) and (4.2), respectively. The table reports the empirical size of testing the joint null hypothesis $\mathbb{H}_{0}:\beta_{1}^{*}=\gamma_{1}^{*}=0$ at a 5% nominal significance level. “Size” is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[\text{Wald}_{\mathcal{J}}^{\text{XD}(r)}>\chi_{0.95,2}^{2}\right]$ across $R=2,000$ replications, where $\text{Wald}_{\mathcal{J}}^{\text{XD}(r)}$ is computed based on (2.19) for the $r$-th replication with $\mathcal{J}=\left\{1,p_{x}+1\right\}$, and the critical value $\chi_{0.95,2}^{2}(\approx 5.99)$ is the 95-th percentile of the chi-squared distribution with 2 degrees of freedom. The “Calibrated” and “CV” columns refer to the methods used for choosing the tuning parameters through calibration and cross-validation, respectively.
@@
We now turn to the empirical power of the XDlasso inference. Figure 2 plots power curves for the null hypotheses $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$ under varying true coefficient values. In our analysis, we vary either $\beta_{1}^{*}$ or $\gamma_{1}^{*}$ from 0 to $0.5$. All remaining coefficients are held fixed as specified in (4.4). Across various configurations, XDlasso exhibits increasingly high power against the null hypothesis as the true coefficient moves away from 0 and the sample size $n$ increases. Furthermore, Figure 3 compares the power for testing $\beta_{1}^{*}$ and $\gamma_{1}^{*}$, which reveals that the power associated with a unit root regressor surpasses that of a stationary regressor. This observation provides finite sample evidence to support the theoretical results in Theorem 2. Specifically, it corroborates the faster convergence of standard error for unit root regressors compared with stationary regressors, thereby inducing higher power in the hypothesis testing.
@@
Notes: The left and right panels correspond to DGPs (4.1) and (4.2), respectively. In each subplot, the first row depicts the empirical power function for $\beta_{1}^{*}$, associated with a nonstationary regressor, across various $\left(p_{x},p_{z}\right)$ configurations, while the second row pertains to $\gamma_{1}^{*}$, associated with a stationary regressor. The empirical power is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (2.15) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}\left(\approx 1.96\right)$ is the 97.5-th percentile of the standard normal distribution.
@@
Notes: This figure plots the power curves under $n=600.$ The first and second rows correspond to DGPs (4.2) and (4.1), respectively. In each subplot, blue lines represent the empirical power function for $\beta_{1}^{*}$, associated with a nonstationary regressor, while red lines represent that for $\gamma_{1}^{*}$, associated with a stationary regressor. Empirical power is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (2.15) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}\left(\approx 1.96\right)$ is the 97.5-th percentile of the standard normal distribution.
@@ 5 Empirical Applications
In recent years, high dimensional macroeconomic data have been extensively used to forecast key variables of interest; see smeekes2018macroeconomic, medeiros2021forecasting, and giannone2021economic, for example. Researchers have primarily focused on the point estimation of forecast. There has been limited empirical exploration of statistical inference on the predictive power of specific predictors, due to a lack of suitable toolkits. This section showcases two empirical applications using our proposed XDlasso inference method. We utilize the monthly data of 112 U.S. macroeconomic variables spanning from January 1960 to April 2025, sourced from the FRED-MD dataset by mccracken2016fred.